OpenStudy (zeesbrat3):

Determine the slope of the graph of x4 = ln(xy) at the point (1, e).

OpenStudy (freckles):

find the slope by first differentiating both sides of the given equation

OpenStudy (zeesbrat3):

But hot can you find the derivative of e?

OpenStudy (zeesbrat3):

@freckles

OpenStudy (freckles):

why do you want the derivative of e? by the way e is a constant and derivative of e is 0. But why do you need this information?

OpenStudy (freckles):

\[\frac{d}{dx}x^4=? \\ \frac{d}{dx}\ln(xy)=?\] the first can be found by using power rule second can be found using chain rule and product rule

OpenStudy (zeesbrat3):

Sorry, I'm just confused as to how to solve for the answer.. This question is on a practice test so I need to learn how to do it for the actual test.

OpenStudy (zeesbrat3):

So the first would be \[4x^3\] And the second would be \[\frac{ d }{ dx } \ln x + \ln y\]

OpenStudy (freckles):

oh okay so you used product rule for log but you still need to differentiate that sum: \[\frac{d}{dx}(\ln(x)+\ln(y)) \\ =\frac{d}{dx}\ln(x)+\frac{d}{dx}\ln(y)\] =?

OpenStudy (zeesbrat3):

\[\frac{ y }{ x }\]?

OpenStudy (freckles):

do you know that: where u is a function of x that: \[\frac{d}{dx}\ln(u)= \frac{u'}{u}\]

OpenStudy (freckles):

\[\frac{d}{dx}\ln(x)=\frac{(x)'}{x}=\frac{1}{x} \\ \frac{d}{dx}\ln(y)=\frac{y'}{y}\]

OpenStudy (freckles):

\[x^4=\ln(xy) \\ \text{ after differentiating both sides we have } \\ 4x^3=\frac{1}{x}+\frac{y'}{y} \] now you can replace x with 1 and y with e and solve for y' (the slope at (1,e))

OpenStudy (zeesbrat3):

So \[4x^3 = 1 + \frac{ y' }{ e }\]

OpenStudy (freckles):

close you need to replace all the x's with 1

OpenStudy (zeesbrat3):

Ooo whoops! \[4 = 1 + \frac{ y' }{ }\]

OpenStudy (zeesbrat3):

theres supposed to be an e under the y prime

OpenStudy (freckles):

\[4=1+\frac{y'}{e}\] yep now just solve for y' (this will be the slope at (1,e) since we replaced x with 1 and y with e)

OpenStudy (zeesbrat3):

So \[3e = y' ?\]

OpenStudy (freckles):

yep

OpenStudy (zeesbrat3):

Thank you! Would you mind helping me with one more? Logs through me for a loop

OpenStudy (freckles):

sure

OpenStudy (zeesbrat3):

Thanks! \[\frac{ d }{ dx }[x^3e^x] = x^2e^x(3x+2)\]

OpenStudy (zeesbrat3):

I have to find whether that is true or not

OpenStudy (freckles):

well do you know product rule?

OpenStudy (freckles):

x^3e^x is a product of x^3 and e^x

OpenStudy (freckles):

\[\frac{d}{dx}(x^3e^x)=e^x \frac{d}{dx}x^3+x^3 \frac{d}{dx}e^x\] what are the following: \[\frac{d}{dx}x^3=? \\ \frac{d}{dx}e^x=?\]

OpenStudy (zeesbrat3):

Right. So the first is \[3x^2\] and the second is 0?

OpenStudy (freckles):

the second is not 0 e^x is not a constant

OpenStudy (freckles):

for example e^2 and e^3 aren't the same so there is no way e^x remains the same

OpenStudy (freckles):

the derivative of e^x is e^x

OpenStudy (zeesbrat3):

Because it is a constant?

OpenStudy (freckles):

I thought I just said it wasn't a constant and gave a reason why it wasn't a constant. e is a constant e^x is not a constant (it varies) (Example: e^2 is not the same as e^3)